Write a system of linear equations, and then use the system to solve the problem.
A cup of uncooked rice contains 15 g of protein and 810 calories. A cup of uncooked soybeans contains 22.5 g of protein and 270 calories. How many cups each should be used for a meal containing 9.5 g of protein and 324 calories?

Answer by scott8148(5882)

let r equal cups of rice and s equal cups of soybeans

r(15)+s(22.5)=9.5 ... r(810)+s(270)=324 ... multiply first equation by -12 and add the second equation

-180r-270s+810r+270s=-114+324 ... 630r=210 ... r=1/3 ... from first equation (1/3)(15)+s(22.5)=9.5 ... s(22.5)=4.5 ...s=1/5